A small tourism company has two old tour buses that frequently break down.
The probability that the first bus is working is 0.60, the probability that the second bus is working is 0.58, and the probability that both are working is 0.45.
What is the probability that neither bus is working?
Express your answer as a probability rounded to two decimal places.

Answer by ikleyn(53763)

.
The probability that at least one bus is working (either bus is working) is


    P(A U B) = P(A) + P(B) - P(AnB) = 0.6 + 0.58 - 0.45 = 0.73.



The probability that NEITHER bus is working is the complement to it


    P(neitherA norB) = 1 - 0.73 = 0.27.

Solved.